Calculate the enthalpies of methanol if the enthalpies of formation of CO2,H2O and CH3OH are respectively -393.5kj/mol,-285.8kj/mol and -239.0kj/mol.
1
Expert's answer
2020-04-06T12:00:21-0400

"\Delta_сH^0=\displaystyle\sum_{i=1}v_i\Delta_fH^0(B_i)-\displaystyle\sum_{i=1}v_i\Delta_fH^0(A_i)"

Where B are products and A are initial reagents

Equation for combustion of methanol:

CH3OH + 1.5O2 = CO2 + 2H2O

Then

"\Delta_сH^0(CH_3OH)=\Delta_fH^0(CO_2)+2\Delta_fH^0(H_2O)" "-1.5\Delta_fH^0(O_2)-\Delta_fH^0(CH_3OH)"

For all simple substances (including O2) under normal conditions enthalpies of formation are taken as 0. So "\Delta_fH^0(O_2) = 0" and

"\Delta_сH^0(CH_3OH)= -393.5 + 2*(-285.8)""-1.5*0-(-239.0)=-726.1" kj/mol